# Predicting Case Suffixes With Activity Start and End Times: A Sweep-Line Based Approach

### v1

@@ 3.1.4 Sequence Encoding and Construction:
Our approach uses two feature types: intra-case features from individual cases and inter-case features from all active cases. These are encoded into feature vectors for each $\mathcal{P}$, combining both contexts to form input sequences.
@@
To represent the input data for the LSTM, we encode both intra-case and inter-case features in a structured and consistent manner. Intra-case features include attributes specific to each event in a case, such as the activity label, inter-start time, and processing time. Activity labels, being categorical, are embedded using trainable embedding layers rather than one-hot encoding. These embeddings are learned jointly with the LSTM network during supervised training, allowing the model to capture task-specific semantic relationships between activities and efficiently handle a larger set of categories, following the approach in [4]. The temporal features, inter-start time and processing time are treated as continuous variables and normalized to the [0,1] range using min-max normalization based on the training data, as illustrated in Table 2.
@@
Inter-case features, such as WIP, resource utilization, and demand rate ($\lambda$), capture the broader system context. These features are updated at each time step and normalized using min-max scaling for stability. Each feature is aligned with its corresponding event to maintain temporal coherence and provide accurate context at every step.
@@
To construct input sequences and target labels for training, we adopted a fixed-size n-gram approach for each $\mathcal{P}$ in $\mathcal{L}$, as outlined in previous studies [4, 19]. Since variable-length prefixes cannot be directly used as inputs, the n-gram method enables standardization of the temporal dimensionality of the input data hence, capturing sub-sequence patterns of activity instances in $\mathcal{P}$. Each n-gram sequence incorporates both intra-case features (e.g., activity, inter-start time, processing time) and inter-case features (e.g., WIP, resource utilization, $\lambda$), ensuring the model captures both case-specific and system-wide process dynamics. Sequences are generated independently for each $\mathcal{P}$, maintaining consistency across all instances, as shown in Table. 2.
3.1.5 Model Architecture Design and Training:
@@
Existing SM architectures predicts all aspects of the next activity instance in a trace prefix $\mathcal{P}$, i.e., the next activity, its inter-start time, and processing time jointly. However, this joint prediction approach lacks flexibility in scenarios where only partial predictions are required. For example, if the activity and its start time are already known, as is the case for ongoing activities, it is often enough to predict only the processing time. In SM-based architectures, this is difficult because all aspects of the next activity instance are predicted together, and the model cannot easily make just one specific prediction. To address this, we propose a BiLSTM-based Multi-model Predictive Learning Architecture (MM) [7], which leverages bidirectional LSTMs to predict the next activity, inter-start time, and processing time for a trace prefix $\mathcal{P}$. BiLSTM analyzes event sequences bidirectionally, capturing activity dependencies, temporal variations, and overlaps.
@@
MM combines categorical and continuous inputs using pre-trained embeddings, concatenated features, and BiLSTM layers to model sequential dependencies. It follows a modular design with three independent BiLSTM predictors: $\alpha$ for next activity, $\beta$ for inter-start time ($\delta^{\text{start}}_{c,k}$), and $\gamma$ for processing time ($\delta^{\text{proc}}_{c,k}$), as shown in Fig. 1. Each model is trained separately on $\mathcal{L}_{train}$ to capture distinct aspects of the process.
@@
To train the next activity predictor $\alpha$, we extract $\mathcal{P}$’s from $\mathcal{L}_{train}$ and construct inter and intra-case feature vectors using Alg. 1. These features are encoded into input sequences with target labels to train $\alpha$. For the inter-start time predictor $\beta$, we use a similar feature set, but the model additionally takes the actual next activity $a_{c,k+1}$ from $\text{$\mathcal{P}$}_{k+1}$ as input to learn accurate temporal dependencies. Fixed-length input sequences and corresponding labels are generated for training. The processing time predictor $\gamma$ also uses inter- and intra-case features, but focuses on estimating the duration of the next activity $a_{c,k+1}$, incorporating both $a_{c,k+1}$ and its inter-start time $\delta^{\text{start}}_{c,k+1}$. This enables $\gamma$ to capture temporal patterns and case-specific variations for accurate processing time prediction.
@@ 3.2.1 Input:
Each $\text{$\mathcal{P}$}{}_{\emptyset}$ is linked to a case identifier and contains the activity name, start timestamp, and if available end timestamp for each instance in $\sigma_{c}{}$, as illustrated in Table 1. These prefixes collectively form the test log ($L^{\prime}$) (line 1). Providing $\text{$\mathcal{P}$}{}_{\emptyset}$ as input to MM offers the historical context required to predict the next activity and its expected start and end timestamps for each ongoing case.
@@ 7  Let $P_{\emptyset}\leftarrow$ $\texttt{Prefix}(L^{\prime},t_{\text{sweep}})$;
10     foreach $r=(c,a,T^{\text{start}},T^{\text{end}})\in P_{c}$ where $T^{\text{end}}=\emptyset$ do
11        Enhance $P_{c}$ with intra/inter-case features using Alg. 1;
@@ 13        $T^{\text{end}}\leftarrow T^{\text{start}}+\delta^{\text{proc}}$;
14        $r\leftarrow(c,a,\delta^{\text{proc}},T^{\text{start}},T^{\text{end}})$;
@@
19     Enhance $P_{c}$ with intra/inter-case features using Alg. 1;
20
@@ 36     $t_{\text{sweep}}\leftarrow\min(t)$
To predict suffixes of all ongoing cases in lock step as of $t_{\text{cutoff}}$, the Alg. 2 begins by initializing the $t_{sweep}$ at $t_{\text{cutoff}}$ (line 2), which serves as the starting point for the sweep-line algorithm to begin completing the incomplete trace prefixes ($\text{$\mathcal{P}$}{}_{\emptyset}$). MM integrates three models $\alpha,\beta,\gamma$ in to a sweep-line based simulation algorithm to sequentially predict the next activity, its inter-start time and its processing time for a given incomplete trace prefix $\text{$\mathcal{P}$}{}_{\emptyset}$.
While $t_{sweep}$ is less than the maximum start timestamp in $\mathcal{L}^{\prime}$, the algorithm extracts incomplete trace prefixes ($\text{$\mathcal{P}$}_{\emptyset}$) as of $t_{sweep}$ (line 4). Each $\text{$\mathcal{P}$}_{\emptyset}$ includes traces where the last activity has started but not yet finished. Lines (5-10) handle such cases by first completing these ongoing activities. For each prefix $\text{$\mathcal{P}$}_{c,k-1}$, inter- and intra-case features are extracted using Alg. 1 (line 7) and passed to the prediction model $\gamma$ to estimate the remaining processing time (line 8). The end timestamp is then computed by adding the predicted $\delta^{\text{proc}}_{c,k}{}$ to the start time of $\text{a}_{c,k}$ (line 9). To support concurrent activities, the algorithm applies this prediction to all activity instances in the prefix $P_{c}$ with missing end timestamps, rather than only the most recent one, ensuring consistent suffix prediction in the presence of parallelism.
With complete activity instances, $\text{$\mathcal{P}$}_{c}$ is enhanced with inter & intra-case features using Alg. 1 (line 12). For each enhanced $\text{$\mathcal{P}$}_{c}$ (lines 13-15), MM sequentially applies its models, where $\alpha$ predicts the next activity instance $\text{act}_{c,k+1}$ given trace $\text{$\mathcal{P}$}_{c}$. $\beta$ predicts its inter-start time $\delta^{\text{start}}_{c,k+1}$ given $\text{a}_{c,k+1}$ and $\text{$\mathcal{P}$}_{c}$ and $\gamma$ predicts its processing time given $\text{a}_{c,k+1}$, $\delta^{\text{start}}_{c,k+1}$, and $\text{$\mathcal{P}$}_{c}$.
@@ 4 Evaluation
Previous methods use Single Model Architectures (SM), where a single model is trained to predict both the next activity and its timestamp. In contrast, as discussed in Sec. 3, our MM approach applies separate models sequentially to predict the next activity, its inter-start time, and its processing time. This modular design offers greater flexibility, especially when the activity and start time are already known at a given $t_{sweep}$, enabling more accurate processing time predictions. However, it remains to be seen whether this modularity affects predictive performance. The evaluation below addresses the question: How does MM compare to SM in predicting the suffix of ongoing cases in a prefix log?
@@
    EQ1: In terms of control flow prediction? We hypothesize that MM’s flexibility – enabling the integration of the most suitable control flow prediction model for each event log through targeted training – enables it to outperform SM-based baselines. We also expect that the inter-case features (resource utilization, demand, and capacity) will not enhance control-flow accuracy, as they capture workload properties rather than execution order.
@@
    EQ2: In terms of inter-start time prediction? Since inter-start times (i.e., time from the start of the current activity to the start of the next activity) are influenced by resource utilization, demand, and capacity, we expect that incorporating inter-case features into our sweep-line approach will improve their prediction. Inter-case features provide needed information on workload and availability, which are determinants of waiting time. Furthermore, we hypothesize that better control flow prediction will further enhance inter-start time estimation. Since activities have varying waiting times, a better next-activity predictor improves inter-start time prediction and hence start timestamp estimation.
  3. 3.
    EQ3: In terms of processing time prediction? Unlike inter-start times, processing times are less dependent on workload factors such as resource utilization, demand and capacity. Instead, they are primarily determined by the nature of the activity itself. Consequently, we do not expect inter-case features to improve processing time prediction. However, we anticipate that a more accurate next-activity prediction could improve processing time estimates. Since activities have distinct processing times, accurate next-activity prediction enhances processing time estimation.

### v2

@@ 3.1.5 Feature Vector Construction
Our approach uses three feature types: intra-case features from individual cases, resource contention features from all active cases and resource availability features extracted from resource calendars. These are encoded into feature vectors for each $\mathcal{P}$, combining all contexts to form input sequences.
@@
To represent the input data for the LSTM, we encode intra-case, resource contention (${R}_{\text{cont}}$) and resource availability (${R}_{\text{avail}}$) features in a structured and consistent manner. Intra-case features include attributes specific to each event in a case, such as the activity label, inter-start time, and processing time. Activity labels, being categorical, are embedded using trainable embedding layers rather than one-hot encoding. These embeddings are learned jointly with the LSTM network during supervised training, allowing the model to capture task-specific semantic relationships between activities and efficiently handle a larger set of categories, following the approach in [5]. The temporal features, inter-start time and processing time are treated as continuous variables and normalized to the [0,1] range using min-max normalization based on the training data, as illustrated in Table 2.
@@
Resource contention features (i.e. WIP, resource utilization & demand rate ($\lambda$)) capture the broader system context, while resource availability features (i.e. ${R}_{\text{avail}}$) represent the current availability of resources. These features are updated at each time step and normalized using min-max scaling to ensure standardization of data. Each feature is aligned with its corresponding event to maintain temporal coherence and provide accurate contextual information at every step.
@@
To construct input sequences and target labels for training, we adopted a fixed-size n-gram approach for each $\mathcal{P}$ in $\mathcal{L}$, as outlined in previous studies [5, 32]. Since variable-length prefixes cannot be directly used as inputs, the n-gram method enables standardization of the temporal dimensionality of the input data hence, capturing sub-sequence patterns of activity instances in $\mathcal{P}$. Each n-gram sequence incorporates intra-case features (e.g., activity, inter-start time, processing time), ${R}_{\text{cont}}$ features (e.g., WIP, resource utilization, $\lambda$) and ${R}_{\text{avail}}$ features ensuring the model captures both case-specific and system-wide process dynamics. Sequences are generated independently for each $\mathcal{P}$, maintaining consistency across all instances, as shown in Table. 2.
3.1.6 Model Architecture Design and Training
@@
Existing SM architectures predicts all aspects of the next activity instance in a trace prefix $\mathcal{P}$, i.e., the next activity, its inter-start time, and processing time jointly. However, this joint prediction approach lacks flexibility in scenarios where only partial predictions are required. For example, if the activity and its start time are already known, as is the case for ongoing activities, it is often enough to predict only the processing time. In SM-based architectures, this is difficult because all aspects of the next activity instance are predicted together, and the model cannot easily make just one specific prediction. To address this, we propose a BiLSTM-based Multi-model Predictive Learning Architecture ($M$) [12], which leverages bidirectional LSTMs to predict the next activity, inter-start time, and processing time for a trace prefix $\mathcal{P}$. BiLSTM analyzes event sequences bidirectionally, capturing activity dependencies, temporal variations, and overlaps.
@@
$M$ combines categorical and continuous inputs using pre-trained embeddings, concatenated features, and BiLSTM layers to model sequential dependencies. It follows a modular design with three independent BiLSTM predictors: $\alpha$ for next activity, $\beta$ for inter-start time ($\delta^{\text{start}}_{c,k}$), and $\gamma$ for processing time ($\delta^{\text{proc}}_{c,k}$), as shown in Fig. 1. Each model is trained separately on $\mathcal{L}_{train}$ to capture distinct aspects of the process.
@@
To train the next activity predictor $\alpha$, we extract $\mathcal{P}$’s from $\mathcal{L}_{train}$ and construct intra-case, ${R}_{\text{cont}}$ and ${R}_{\text{avail}}$ feature vectors using Alg. 1 to capture both process dynamics and resource context. These features are encoded into input sequences with target labels to train $\alpha$. For the inter-start time predictor $\beta$, we use a similar feature set, but the model additionally takes the actual next activity $a_{c,k+1}$ from $\text{$\mathcal{P}$}_{k+1}$ as input to learn accurate temporal dependencies. Fixed-length input sequences and corresponding labels are generated for training. The processing time predictor $\gamma$ also uses similar features, but focuses on estimating the duration of the next activity $a_{c,k+1}$, incorporating both $a_{c,k+1}$ and its inter-start time $\delta^{\text{start}}_{c,k+1}$. This enables $\gamma$ to capture temporal patterns and case-specific variations for accurate processing time prediction.
3.1.7 Output
The offline phase produces three independently trained BiLSTM predictors. Given a trace prefix $\mathcal{P}$, the activity predictor $\alpha$ outputs a probability distribution over possible next activities. Conditioned on $\mathcal{P}$ and the next activity $a_{c,k+1}$, the inter-start time predictor $\beta$ estimates $\delta^{\text{start}}_{c,k+1}$, while the processing time predictor $\gamma$ estimates $\delta^{\text{proc}}_{c,k+1}$ based on $\mathcal{P}$, $a_{c,k+1}$, and $\delta^{\text{start}}_{c,k+1}$. These predictors are selectively used in the online phase depending on the available runtime information.
@@ 3.2.1 Input
Each $\text{$\mathcal{P}$}{}_{\emptyset}$ is linked to a case identifier and contains the activity name, start timestamp, and if available end timestamp for each instance in $\sigma_{c}{}$, as illustrated in Table 1. These prefixes collectively form the test log ($L^{\prime}$) (line 1). Providing $\text{$\mathcal{P}$}{}_{\emptyset}$ as input to $M$ offers the historical context required to predict the next activity and its expected start and end timestamps for each ongoing case.
To predict suffixes of all ongoing cases in lock step as of $t_{\text{cutoff}}$, the Alg. 2 begins by initializing the $t_{sweep}$ at $t_{\text{cutoff}}$ (line 2), which serves as the starting point for the sweep-line algorithm to begin completing the incomplete trace prefixes ($\text{$\mathcal{P}$}{}_{\emptyset}$). $M$ integrates three models $\alpha,\beta,\gamma$ in to a sweep-line based simulation algorithm to sequentially predict the next activity, its inter-start time and its processing time for a given incomplete trace prefix $\text{$\mathcal{P}$}{}_{\emptyset}$.
While $t_{sweep}$ is less than the maximum start timestamp in $\mathcal{L}^{\prime}$, the algorithm extracts incomplete trace prefixes ($\text{$\mathcal{P}$}_{\emptyset}$) as of $t_{sweep}$ (line 4). Each $\text{$\mathcal{P}$}_{\emptyset}$ includes traces where the last activity has started but not yet finished. Lines (5-10) handle such cases by first completing these ongoing activities. For each prefix $\text{$\mathcal{P}$}_{c,k-1}$, intra-case, ${R}_{\text{cont}}$ & ${R}_{\text{avail}}$ features are extracted using Alg. 1 (line 7) and passed to the prediction model $\gamma$ to estimate the remaining processing time (line 8). The end timestamp is then computed by adding the predicted $\delta^{\text{proc}}_{c,k}{}$ to the start time of $\text{a}_{c,k}$ (line 9). To support concurrent activities, the algorithm applies this prediction to all activity instances in the prefix $P_{c}$ with missing end timestamps, rather than only the most recent one, ensuring consistent suffix prediction in the presence of parallelism.
With complete activity instances, $\text{$\mathcal{P}$}_{c}$ is enhanced with intra-case, ${R}_{\text{cont}}$ & ${R}_{\text{avail}}$ features using Alg. 1 (line 12). For each enhanced $\text{$\mathcal{P}$}_{c}$ (lines 13-15), $M$ sequentially applies its models, where $\alpha$ predicts the next activity instance $\text{act}_{c,k+1}$ given trace $\text{$\mathcal{P}$}_{c}$. $\beta$ predicts its inter-start time $\delta^{\text{start}}_{c,k+1}$ given $\text{a}_{c,k+1}$ and $\text{$\mathcal{P}$}_{c}$ and $\gamma$ predicts its processing time given $\text{a}_{c,k+1}$, $\delta^{\text{start}}_{c,k+1}$, and $\text{$\mathcal{P}$}_{c}$.
3.2.2 Output
The output of the online phase is a completed activity instance log $L^{\prime}$, where each incomplete trace prefix observed at $t_{\text{cutoff}}$ is extended with a predicted case suffix using the models $\alpha$, $\beta$, and $\gamma$ trained during the offline phase. The predicted suffixes consist of a sequence of future activity instances, each associated with a predicted activity label and corresponding temporal information. Resource prediction for activity instances in the predicted suffixes is not performed, as prior work [5] indicates limited accuracy at this level of granularity; instead, higher-level abstractions such as roles have been shown to be more informative, although role prediction is beyond the scope of this work. Overall, the resulting log $L^{\prime}$ represents a complete and temporally consistent extension of the test log and is directly comparable to a fully observed activity instance log.
@@ 7   Let $P_{\emptyset}\leftarrow$ $\texttt{Prefix}(L^{\prime},t_{\text{sweep}})$;
10    foreach $e=(c,a,T^{\text{start}},T^{\text{end}})\in P_{c}$ where $T^{\text{end}}=\emptyset$ do
11         Enhance $P_{c}$ with intra, ${R}_{\text{cont}}$ & ${R}_{\text{avail}}$ features using Alg. 1;
@@ 13       $T^{\text{end}}\leftarrow T^{\text{start}}+\delta^{\text{proc}}$;
14       $e\leftarrow(c,a,\delta^{\text{proc}},T^{\text{start}},T^{\text{end}})$;
@@
21      Enhance $P_{c}$ with intra, ${R}_{\text{cont}}$ & ${R}_{\text{avail}}$ features using Alg. 1;
22
@@ 4 Evaluation
Previous methods use Single Model Architectures (SM), where a single model is trained to predict both the next activity and its timestamp. In contrast, as discussed in Sec. 3, our $M$ approach applies separate models sequentially to predict the next activity, its inter-start time, and its processing time. This modular design offers greater flexibility, especially when the activity and start time are already known at a given $t_{sweep}$, which enables more accurate processing time predictions. However, it remains to be seen whether this modularity affects predictive performance. The evaluation below addresses the following questions:
@@
    EQ1: How does $M$ compare to SM in predicting the suffix of ongoing cases in a prefix log?
    1. (a)
      EQ1(a) In terms of control flow prediction? We hypothesize that $M$’s flexibility – enabling the integration of the most suitable control flow prediction model for each event log through targeted training – enables it to outperform SM-based baselines. We also expect that the resource contention features (resource utilization, demand, and capacity) will not enhance control-flow accuracy, as they capture workload properties rather than execution order.
    2. (b)
      EQ1(b) In terms of inter-start time prediction? Since inter-start times (i.e., time from the start of the current activity to the start of the next activity) are influenced by resource utilization, demand, and capacity, we expect that incorporating resource contention features into our sweep-line approach will improve their prediction. Resource contention features provide needed information on workload and availability, which are determinants of waiting time. Furthermore, we hypothesize that better control flow prediction will further enhance inter-start time estimation. Since activities have varying waiting times, a better next-activity predictor improves inter-start time prediction and hence start timestamp estimation.
    3. (c)
      EQ1(c) In terms of processing time prediction? Unlike inter-start times, processing times are less dependent on workload factors such as resource utilization, demand and capacity. Instead, they are primarily determined by the nature of the activity itself. Consequently, we do not expect resource contention features to improve processing time prediction. However, we anticipate that a more accurate next-activity prediction could improve processing time estimates. Since activities have distinct processing times, accurate next-activity prediction enhances processing time estimation.
@@
    EQ2: How does the inclusion of resource availability features influence $M$’s predictive performance? While EQ1 compares the predictive power of $M$ and SM, EQ2 compares $M$ and MMavail (where MMavail represents the model with resource-availability features), isolating the impact of resource-availability features. Specifically, we assess whether incorporating information on resource availability enable improved predictions across control flow, inter-start time, and processing time tasks. We hypothesize that incorporating resource availability features can improve predictive performance across all three tasks. When information about which resources are available at a given time is included, the model can better infer which activities are more likely to occur next, thereby enhancing control flow prediction. Moreover, since the start of the next activity often depends on when a suitable resource becomes available, these features are expected to substantially improve inter-start time prediction. Finally, as both the control-flow and inter-start times are predicted more accurately, this most likely yield more precise processing time predictions.
4.1 Experiment Setup
In SM-based experiments, activity instances that start and finish before the cutoff are excluded from the test set, as SM predicts the next activity, inter-start time, and processing time together but cannot handle ongoing instances. In contrast, $M$ includes ongoing instances since it can predict their processing time given its activity and start timestamp.
Experiments were conducted on a desktop with an NVIDIA RTX 3090 GPU, Intel Core i9 CPU, and 64 GB RAM. Training the three $M$ models ($\alpha$, $\beta$, $\gamma$) takes 2.5 hours for smaller logs (e.g., ACR, CFS) and 5-6 hours for larger ones (e.g., CVS, BPI2017W). During testing, the sweep-line approach predicts suffixes of ongoing cases at runtime. For BPI2017W, the approach generated 15,000 predicted events in 100 seconds at 150 events per second.
4.1.1 Next Activity Prediction in the $M$ Architecture
A core component of the $M$ architecture is a model $\alpha$ that predicts the next activity in a case. Using an LSTM-based approach, the model generates multiple possible next activities, each assigned a probability score. A sampling method then selects the next activity. We evaluated $M$’s accuracy with three sampling methods: Argmax, which selects the most probable activity [32]; Random Choice, which randomly selects an activity based on the probability distribution [5]; and Daemon Action, a heuristic-based method [1]. All three methods performed similarly, but Daemon Action produced the most consistent results across datasets. Therefore, we selected Daemon Action as the preferred sampling method for model $\alpha$.
4.1.2 Hyperparameter Optimization
We applied hyperparameter optimization within the sets of values in Table 3 for selecting the most suitable three $M$ models ($\alpha$ for next activity, $\beta$ for inter-start time, $\gamma$ for processing time). Batch size is optimized for computational efficiency and gradient stability, while normalization methods like log normalization and max scaling improve convergence. The n-gram size (N_size) ensures meaningful historical dependencies, and different LSTM layer sizes (L_size) balance model complexity and training efficiency. Activation functions (Selu, Tanh) help maintain stable gradients, and optimizers (Nadam, Adam, SGD, Adagrad) are chosen for optimal convergence. The training set is split (80% training, 20% validation), and $M$ is trained with diverse hyperparameter combinations. We optimized hyperparameters using a random search over 50 iterations, which provides a practical balance between search space coverage and computational efficiency. Each model was trained for up to 200 epochs with early stopping(patience = 10) to prevent overfitting and reduce unnecessary training time. These settings are consistent with prior studies [5, 1] and were empirically validated to ensure stable convergence.
Table 3: Hyperparameter settings for LSTM model
Parameter Explanation Search Space
Batch size Number of samples per update {32, 64, 128}
Normalization Input scaling method {lognorm, max}
Epochs Training iterations 200
N_size Size of n-gram window {5, 10, 15, 20, 25}
L_size LSTM layer size {50, 100, 150}
Activation Hidden layer activation {selu, tanh}
Optimizer Optimization algorithm {Nadam, Adam, SGD, Adagrad}
During optimization, we evaluated performance using two metrics: Mean Absolute Error (MAE)  [^6] MAE is defined as $\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|$ for $\delta^{\text{start}}_{c,k}$ and $\delta^{\text{proc}}_{c,k}$ predictions; and Categorical Cross-Entropy Loss [^7] Categorical Cross-Entropy Loss is defined as $\mathcal{L}_{\text{CE}}=-\sum_{i=1}^{n}\sum_{j=1}^{C}y_{ij}\log(\hat{y}_{ij})$ for next-activity prediction.
To evaluate BiLSTM configuration impact, we tested multiple $M$ variants with varying BiLSTM layer sizes and n-gram window sizes (see Table 3). Larger models marginally improved accuracy but increased training time and reduced throughput. The default configuration (100-unit layer size, 10 n-gram size, 64 batch size, 200 epochs, Adam optimizer) strikes a practical balance between accuracy and efficiency, offering an effective trade-off for predicting case suffixes in business processes.
4.1.3 Measures of Goodness
4.1.4 Datasets
We evaluate our approach using nine event logs [^8] All logs, including a mix of large- and small-scale processes, are available at supplementary material https://shorturl.at/PM3aV, selected for their diverse control-flow structures and temporal characteristics. Table 4 reports for each log the number of cases(#Cases), activity instances (#Act-Ins), and unique activities (#Act). (#Act/Case) is the average number of activities per case. CV Len and CV Dur denotes the coefficient of variation in case length and duration respectively [^9] The coefficient of variation (CV) is computed as: $\text{CV}=\frac{\sigma}{\mu}$, where $\sigma$ is the standard deviation and $\mu$ is the mean of case length or duration.. Avg.Dur and Max.Dur shows average and maximum case durations in days.